Five Measures of Solvency or Profitability The balance sheet for Garcon Inc. at the end of...

  1. Five Measures of Solvency or Profitability

    The balance sheet for Garcon Inc. at the end of the current fiscal year indicated the following:

    Bonds payable, 8% $1,000,000
    Preferred $5 stock, $50 par 146,000
    Common stock, $11 par 192,720.00

    Income before income tax was $224,000, and income taxes were $34,200 for the current year. Cash dividends paid on common stock during the current year totaled $43,800. The common stock was selling for $100 per share at the end of the year.

    Determine each of the following. Round answers to one decimal place, except for dollar amounts which should be rounded to the nearest whole cent. Use the rounded answers for subsequent requirements, if required.

    a. Times interest earned ratio times
    b. Earnings per share on common stock $
    c. Price-earnings ratio
    d. Dividends per share of common stock $
    e. Dividend yield %
Answer #1
a. Times interest earned ratio 3.8 times
b. Earnings per share on common stock $    10.00
c. Price-earnings ratio 10
d. Dividends per share of common stock $       2.50
e. Dividend yield 2.5 %

Workings:

a. Times interest earned ratio = EBIT/Interest
EBIT = EBT + Interest = $224000 + $80000 = $304000
Interest = 8% x $1000000 = $80000
Times interest earned ratio = $304000/$80000 = 3.8
b. Earnings per share on common stock = (Net income - Preferred dividend)/Common stock outstanding
Net income = EBT - Income taxes = $224000 - $34200 = $189800
Preferred dividend = Number of preferred stock x $5 = $146000/$50 x $5 = $14600
Common stock outstanding = $192720/$11 = 17520
Earnings per share on common stock = ($189800 - $14600)/17520 = $175200/17520 = $10
c. Price-earnings ratio = Market price per share/Earnings per share = $100/$10 = 10
d. Dividends per share of common stock = Cash dividends paid on common stock/Number of common stock = $43800/17520 = $2.50
e. Dividend yield = Dividend per share/Market price per share = $2.50/$100 = 2.5%